OpenStudy (anonymous):

If the equation x^3 + px + q = 0 has real roots then prove that 27q^2 + 4p^3 < 0.

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (inkyvoyd):

lool I'm sleepy?

OpenStudy (inkyvoyd):

3 real roots?

OpenStudy (anonymous):

YEp.

OpenStudy (inkyvoyd):

Start by reading wikipedia xD

OpenStudy (asnaseer):

there are well known formulas for the solution of a "reduced" cubic equation (which is what you have). see here for example: http://www.trans4mind.com/personal_development/mathematics/polynomials/cubicAlgebra.htm

OpenStudy (anonymous):

No. It must have a more direct solution.

OpenStudy (anonymous):

you study dirivative?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

i think f(x)=x^3+px+q find solution f'(x)=0 is x1 anh x2 f(x1)*f(x2)<0

OpenStudy (anonymous):

Are you getting the req. expression? Cause Im not.

OpenStudy (anonymous):

????

OpenStudy (anonymous):

A cubic \[a x^3+b x^2+c x+d=0 \] has three real roots if the discriminant \[ \Delta =-27 a^2 d^2+18 a b c d-4 a c^3-4 b^3 d+b^2 c^2 > 0 \] in our case a = 1; b = 0; c = p; d = q; then \[\Delta=-4 p^3-27 q^2 \] we are done

OpenStudy (inkyvoyd):

@eliassaab , nice job, but where'd you get the discriminant and the has three real roots from?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Cubic_function

OpenStudy (anonymous):

Damn that formula. Nice work.